### v2

# Multi-student Diffusion Distillation for Better One-step Generators

Abstract
Diffusion models achieve high-quality sample generation at the cost of a lengthy multistep inference procedure. To overcome this, diffusion distillation techniques produce student generators capable of matching or surpassing the teacher in a single step. However, the student model’s inference speed is limited by the size of the teacher architecture, preventing real-time generation for computationally heavy applications. In this work, we introduce Multi-Student Distillation (MSD), a framework to distill a conditional teacher diffusion model into multiple single-step generators. Each student generator is responsible for a subset of the conditioning data, thereby obtaining higher generation quality for the same capacity. MSD trains multiple distilled students, allowing smaller sizes and, therefore, faster inference. Also, MSD offers a lightweight quality boost over single-student distillation with the same architecture. We demonstrate MSD is effective by training multiple same-sized or smaller students on single-step distillation using distribution matching and adversarial distillation techniques. With smaller students, MSD gets competitive results with faster inference for single-step generation. Using $4$ same-sized students, MSD significantly outperforms single-student baseline counterparts and achieves remarkable FID scores for one-step image generation: $1.20$ on ImageNet-64$\times$64 and $8.20$ on zero-shot COCO2014.  [^1] Project page: https://research.nvidia.com/labs/toronto-ai/MSD/
1 Introduction
Diffusion models are the dominant generative model in image, audio, video, 3D assets, protein design, and more [20, 28, 6, 2, 47]. They allow conditioning inputs – such as class labels, text, or images – and achieve high-quality generated outputs. However, their inference process typically requires hundreds of model evaluations – with an often slow and bulky network – for a single sample. This procedure costs millions of dollars per day [66, 15] and prohibits applications requiring rapid synthesis, such as augmented reality. Real-time, low-cost, and high-quality generation will have enormous financial and operational impacts while enabling new usage paradigms.
There has been a flurry of work on diffusion-distillation techniques to address diffusion model’s slow sampling [41, 65, 76]. Inspired by knowledge distillation [19], these methods use the trained diffusion model as a teacher and optimize a student model to match its output in as little as a single step. However, most diffusion distillation methods use the same student and teacher architecture. This prevents real-time generation for applications with bulky networks, such as video synthesis [6]. While reducing model size can reduce inference time, it typically yields worse generation quality, thus presenting a speed-to-quality tradeoff dilemma with existing distillation methods.
Figure 1: We visualize distilling into multiple students, where each student handles a subset of the input condition. At training, students are trained separately with filtered data. At inference, a single student is retrieved for generation, given the corresponding input condition. We show samples in Fig. 2, with methodological details outlined in Fig. 3, and quantitative results in Tables 2 and 2.
We tackle this dilemma with our method, Multi-Student Distillation (MSD), that introduces multiple single-step student generators distilled from the pretrained teacher. Each student is responsible for a subset of conditioning inputs. We determine which student to use during inference and perform a single-model evaluation to generate a high-quality sample. This way, MSD enjoys the benefit of a mixture of experts [24]: it increases the model capacity without incurring more inference cost, thereby effectively pushing the limit of the speed-quality tradeoff.
When distilling into the same-sized students, MSD has the flexibility of conceptually applying to any distillation method for a performance boost. In addition, MSD allows distilling into multiple smaller students for reduced single-step generation time. Using smaller students prevents one from initializing a student from teacher weights, posing an additional technical challenge. We solve this challenge by adding a relatively lightweight score-matching pretraining stage before distillation (Sec. 4.3) and demonstrating its necessity and efficiency via extensive experiments.
We validate our approaches by applying MSD to distill the teacher into $4$ same-sized students using distribution matching and adversarial distillation procedures [76, 75]. The resulting students collectively outperform single-student counterparts, achieving close to state-of-the-art FID scores of $1.20$ on one-step ImageNet-64$\times$64 generation (Tab. 2) and $8.20$ on one-step zero-shot COCO2014 generation (Tab. 2). We also distill the same teacher into $4$ smaller students, which achieve a competitive FID of $2.88$ on ImageNet (Tab. 2), with $42\%$ less parameters per student.
Refer to captionRefer to captionRefer to caption(a) Teacher (multistep)(b) Same-sized student(c) 83% smaller student
Figure 2: Samples on high guidance-scale text-to-image generations from the SD v1.5 teacher and different sized students, with full training details in App. D. The same-sized student has comparable quality to the teacher. The smaller student, trained on a subset of dog-related data, achieves faster generation while still having decent qualities. The same-sized student is trained with DM stage only, whereas the smaller student is trained with TSM and DM stages (see Fig. 3). See additional samples in Fig. 12.
We summarize our contributions below, which include:
  • A new framework, MSD, that upgrades existing single-step diffusion distillation methods (Sec. 4.1) by increasing the effective model capacity without changing inference latency.
  • Demonstrating the effectiveness of MSD by training multiple same sized students using SOTA distillation techniques [76, 75] in Sec. 4.2, surpassing single-student counterparts in FID scores for ImageNet-64$\times$64 (Sec. 5.2) and zero-shot text-to-image generation (Sec. 5.3).
  • A successful scheme to distill multiple smaller single-step students from the teacher model, achieving comparable generation quality with reduced inference time.
2 Related Work
Diffusion Sampling Acceleration. While a line of work aims to accelerate diffusion models via fast numerical solvers for the PF-ODE [39, 40, 84, 25, 35], they usually still require more than $10$ steps. Training-based methods that usually follow the knowledge distillation pipeline can achieve low-step or even one-step generation. Luhman and Luhman [41] first used the diffusion model to generate a noise and image pair dataset that is then used to train a single-step generator. DSNO [83] precomputes the denoising trajectory and uses neural operators to estimate the whole PF-ODE path. Progressive distillation [55, 44] iteratively halves the number of sampling steps required without needing an offline dataset. Rectified Flow [36] and follow-up works [37, 73] straighten the denoising trajectories to allow sampling in fewer steps. Another approach uses self-consistent properties of denoising trajectories to inject additional regularization for distillation [16, 5, 65, 63, 42, 53, 27].
The methods above require the student to follow the teacher’s trajectories. Instead, a recent line of works aims to only match the distribution of the student and teacher output via variational score distillation [76, 75, 56, 70, 43, 88, 45, 86]. The adversarial loss [14], often combined with the above techniques, has been used to enhance the distillation performance further [69, 82, 59, 60, 67, 72, 33, 27, 87]. Although MSD is conceptually compatible and offers a performance boost to all of these distillation methods, in this work, we demonstrate two specific techniques: distribution matching [76] and adversarial distillation [75].
Mixture of experts training and distillation. Mixture of Experts (MoE), first proposed in [24], has found success in training very large-scale neural networks [62, 30, 13, 31, 7]. Distilling a teacher model into multiple students was explored by Hinton et al. [19], and after that, has been further developed for supervised learning [10, 46, 9] and language modeling [71, 29, 89]. Although several works [21, 48, 1] have proposed MoE training schemes for generative adversarial networks, they train the MoE from scratch. This requires carefully tuning the multi-expert adversarial losses. eDiff-I [3] uses different experts in different denoising timesteps for a multi-step diffusion model. A recent work [85] proposes to distill a pretrained diffusion model into an MoE for policy learning, which shares similar motivations with our work. However, to the best of our knowledge, MSD is the first method to distill multi-step teacher diffusion models into multiple one-step students for image generation.
Efficient architectures for diffusion models. In addition to reducing steps, an orthogonal approach aims to accelerate diffusion models with more efficient architectures. A series of works [4, 50, 22] introduces vision transformers to diffusion blocks and trains the diffusion model with new architectures from scratch. Another line of work selectively removes or modifies certain components of a pretrained diffusion model and then either finetunes [26, 32, 78] or re-trains [81] the lightweight diffusion model, from which step-distillation can be further applied [32, 81]. Our approach is orthogonal to these works in two regards: 1) In our method, each student only handles a subset of data, providing a gain in relative capacity. 2) Instead of obtaining a full diffusion model, our method employs a lightweight pretraining stage to obtain a good initialization for single-step distillation. Combining MSD with more efficient architectures is a promising future direction.
3 Preliminary
We introduce the background on diffusion models in Sec. 3.1 and distribution matching distillation (DMD) in Sec. 3.2. We discuss applying adversarial losses to improve distillation in Sec. 3.3.
3.1 Diffusion Models
[…]
3.2 Distribution Matching Distillation
Inspired by Wang et al. [68], a series of works [43, 76, 74, 45] aim to train the single-step distilled student to match the generated distribution of the teacher diffusion model. This is done by minimizing the following reverse KL divergence between teacher and student output distributions, diffused at different noise levels for better support over the ambient space:
$\mathbb{E}_{t}D_{\text{KL}}(p_{t,\text{fake}}\|p_{t,\text{real}})=\mathbb{E}_{\bm{x}_{t}}\left(\log\left(\frac{p_{t,\text{fake}}(\bm{x}_{t})}{p_{t,\text{real}}(\bm{x}_{t})}\right)\right).$ (1)
The training only requires the gradient of Eq. (1), which reads (with a custom weighting $w_{t}$):
$\nabla_{\theta}\mathcal{L}_{\text{KL}}(\theta):=\nabla_{\theta}\mathbb{E}_{t}D_{\text{KL}}\simeq\mathbb{E}_{\bm{z},t,\bm{x}_{t}}[w_{t}\alpha_{t}(\bm{s}_{\text{fake}}(\bm{x}_{t},t)-\bm{s}_{\text{real}}(\bm{x}_{t},t))\nabla_{\theta}G_{\theta}(\bm{z})],$ (2)
where $\bm{z}\sim\mathcal{N}(0,\bm{I}),t\sim\text{Uniform}[T_{\min},T_{\max}]$, and $\bm{x}_{t}\sim q(\bm{x}_{t}|\bm{x})$, the noise injected version of $\bm{x}=G_{\theta}(\bm{z})$ generated by the one-step student. Here, we assume the teacher denoising model accurately approximates the score of the real data, and a “fake” denoising model approximates the score of generated fake data:
$\bm{s}_{\text{real}}(\bm{x}_{t},t)\approx-\frac{\bm{x}_{t}-\alpha_{t}\bm{\mu}_{\text{teacher}}(\bm{x}_{t},t)}{\sigma_{t}^{2}},\ \ \bm{s}_{\text{fake}}(\bm{x}_{t},t)\approx-\frac{\bm{x}_{t}-\alpha_{t}\bm{\mu}_{\text{fake}}(\bm{x}_{t},t)}{\sigma_{t}^{2}}.$ (3)
The “fake” denoising model is trained with the denoising objective with weighting $\lambda_{t}$:
$\mathcal{L}_{\text{denoise}}(\phi)=\mathbb{E}_{\bm{z},t,\bm{x}_{t}}[\lambda_{t}\|\bm{\mu}_{\text{fake}}^{\phi}(\bm{x}_{t},t)-\bm{x}\|_{2}^{2}].$ (4)
The generator and the “fake” denoising model are updated alternatively. To facilitate better convergence of the KL divergence, Distribution Matching Distillation (DMD) and DMD2 [75] used two distinct strategies, both significantly improving the generation performance. DMD proposes to complement the KL loss with a regression loss to encourage mode covering:
$\mathcal{L}_{\text{reg}}(\theta)=\mathbb{E}_{(\bm{z},y)\sim\mathcal{D}_{\text{paired}}}\ell(G_{\theta}(\bm{z}),y),$ (5)
where $\mathcal{D}_{\text{paired}}$ is a dataset of latent-image pairs generated by the teacher model offline, and $\ell$ is the Learned Perceptual Image Patch Similarity (LPIPS) [79]. DMD2 instead applies a two-timescale update rule (TTUR), where they update the “fake” score model for $N$ steps per generator update, allowing more stable convergence. We use distribution matching (DM) to refer to all relevant techniques introduced in this section.
3.3 Enhancing distillation quality with adversarial loss
The adversarial loss, originally proposed by Goodfellow et al. [14], has shown a remarkable capability in diffusion distillation to enhance sharpness and realism in generated images, thus improving generation quality. Specifically, DMD2 [75] proposes adding a minimal discriminator head to the bottleneck layer of the “fake” denoising model $\mu_{\text{fake}}$, which is naturally compatible with DMD’s alternating training scheme and the TTUR. Moreover, they showed that one should first train the model without GAN to convergence, then add the GAN loss and continue training. This yields better terminal performance than training with the GAN loss from the beginning. We use adversarial distribution matching (ADM) to refer to distribution matching with added adversarial loss.
4 Method
In Sec. 4.1, we introduce the general Multi-Student Distillation (MSD) framework. In Sec. 4.2, we show how MSD is applied to distribution matching and adversarial distillation. In Sec. 4.3, we introduce an additional training stage enabling distilling into smaller students.
Figure 3: Three-stage training scheme in Eq. 9. Acronym meanings: TSM: teacher score matching (Eq. 8 & Eq. 9); DM: distribution matching (Eq. 9 & Sec. 3.2); ADM: adversarial distribution matching (Eq. 9 and Sec. 3.3). Stage 1 and Stage 2 are techniques from previous works that help with same-sized students; Stage 0 is our contribution, which is required for smaller students who cannot initialize with teacher weights.
4.1 Distilling into multiple students
We present Multi-Student Distillation (MSD), a general drop-in framework to be combined with any conditional single-step diffusion distillation method that enables a cheap upgrade of model capacity without impairing the inference speed. We first identify the key components of a single-step diffusion distillation framework and then present the modification of MSD.
In the vanilla one-student distillation, we have a pretrained teacher denoising diffusion model $\bm{\mu}_{\text{teacher}}$, a training dataset $\mathcal{D}$, and a distillation method. The distillation yields a single-step generator $G(\bm{z};y\in\mathcal{Y})$ via $G=\text{Distill}(\bm{\mu}_{\text{teacher}};\mathcal{D})$. The obtained generator $G$ maps a random latent $\bm{z}$ and an input condition $y$ into an image. In comparison, in an MSD scheme, we instead distill the teacher into $K$ different one-step generators $\{G_{k}(\bm{z};y\in\mathcal{Y}_{k})\}_{k=1}^{K}$ via
$G_{k}=\text{Distill}(\mu_{\text{teacher}},\mathcal{D}_{k}=F(\mathcal{D},\mathcal{Y}_{k})),\ \ k=1,...,K$ (6)
Specifically, each distilled student $G_{k}$ is specialized in handling a partitioned subset $\mathcal{Y}_{k}$ of the whole input condition set $\mathcal{Y}$. So, it is trained on a subset of the training data $\mathcal{D}_{k}\subset\mathcal{D}$, determined by $\mathcal{Y}_{k}$ via a filtering function $F$. Fig. 1 illustrates this idea.
The partition of $\mathcal{Y}$ into $\{\mathcal{Y}_{k}\}_{k=1}^{K}$ determines the input condition groups each student is responsible for. As a starting point, we make the following three simplifications for choosing a partition:
  • Disjointness: This prevents potential redundant training and redundant usage of model capacity.
  • Equal size: Since students have the same architecture, the partitions $\{\mathcal{Y}_{k}\}_{k=1}^{K}$ should be of equal size and require similar model capacity.
  • Clustering: Conditions within each partition should be more semantically similar than those in other partitions, so networks require less capacity to achieve a set quality on their partition.
The first two conditions can be easily satisfied in practice, while the third is not straightforward. For a class-conditional generation, partitioning by semantically similar and equal-sized classes serves a straightforward strategy, though extending it to text-conditional generation is nontrivial. Another promising strategy uses pretrained embedding layers such as the CLIP [51] embedding layer or the teacher embedding layer. One could find embeddings of the input conditions and then perform clustering on those embeddings, which are fixed-length numerical vectors containing implicit semantic information. We ablate partition strategies in Sec. 5.4.
The data filtering function $F$ determines the training subset data $\mathcal{D}_{k}$ from $\mathcal{Y}_{k}$. For example, a vanilla filtering strategy could set $F(\mathcal{D},\mathcal{Y}_{k})=\mathcal{D}_{k}:=\mathcal{D}_{\mathcal{Y}_{k}}$, where $\mathcal{D}_{\mathcal{Y}_{k}}$ denotes the training data subset $\mathcal{D}$ containing the desired condition $\mathcal{Y}_{k}$. Empirically, we found that this filtering works in most cases, although sometimes a different approach is justified, as demonstrated in Sec. 4.2.
4.2 MSD with distribution matching
As a concrete example, we demonstrate the MSD framework using distribution matching (DM) and adversarial distillation techniques. Inspired by the two-stage framework in [75], each of our students is trained with a distribution matching scheme at the first stage and finetuned with an additional adversarial loss at the second stage (adversarial distribution matching, or ADM):
$G_{k}^{(1)}$ $=\text{Distill}_{\text{DM}}\left(\bm{\mu}_{\text{teacher}},F_{\text{DM}}(\mathcal{D}_{\text{DM}},\mathcal{Y}_{k})\right),\ \ k=1,...,K,$ (7)
$G_{k}^{(2)}$ $=\text{Distill}_{\text{ADM}}\left(\bm{\mu}_{\text{teacher}},F_{\text{ADM}}(\mathcal{D}_{\text{ADM}},\mathcal{Y}_{k});G_{k}^{(1)}\right),\ \ k=1,...,K,$
where we recall that $\bm{\mu}_{\text{teacher}}$ is the teacher diffusion model, $G_{k}^{(i)}$ is the $k$-th student generator at the $i$-th stage, $F$ is the data filtering function, $\mathcal{D}$ is the training data, and $\mathcal{Y}_{k}$ is the set of labels that student $k$ is responsible of. The first stage $\text{Distill}_{\text{DM}}$ uses distribution matching with either a complemented regression loss or the TTUR, with details in Sec. 3.2. These two methods achieve optimal training efficiency among other best-performing single-step distillation methods [70, 88, 27] without an adversarial loss, with a detailed comparison in App. B.3. The second stage $\text{Distill}_{\text{ADM}}$ adds an additional adversarial loss (detailed in Sec. 3.3), which introduces minimal additional computational overhead and allows resuming from the first stage checkpoint.
Designing the training data
From Sec. 3, the data required for DM and ADM are $\mathcal{D}_{\text{DM}}=(\mathcal{D}_{\text{paired}},\mathcal{C})$ and $\mathcal{D}_{\text{ADM}}=(\mathcal{D}_{\text{real}},\mathcal{C})$, where $\mathcal{D}_{\text{paired}},\mathcal{D}_{\text{real}},\mathcal{C}$ represents generated paired data, real data and separate conditional input, respectively. We now discuss choices for the filtering function.
For the first stage data filtering $F_{\text{DM}}$, we propose $F_{\text{DM}}(\mathcal{D}_{\text{DM}},\mathcal{Y}_{k})=(\mathcal{D}_{\text{paired}},\mathcal{C}_{\mathcal{Y}_{k}})$, where $\mathcal{C}_{\mathcal{Y}_{k}}$ denotes the subset of condition inputs $\mathcal{C}$ that contains $\mathcal{Y}_{k}$. In other words, we sample all input conditions only on the desired partition for the KL loss but use the whole paired dataset for the regression loss. This special filtering is based on the observation that the size of $\mathcal{D}_{\text{paired}}$ critically affects the terminal performance of DMD distillation: using fewer pairs causes mode collapse, whereas using more pairs challenges the model capacity. Naïvely filtering paired datasets by partition reduces the paired dataset size for each student and leads to worse performance, as in our ablation in App. B.2. Instead of generating more paired data to mitigate this imbalance, we simply reuse the original paired dataset for the regression loss. This is remarkably effective, which we hypothesize is because paired data from other input conditions provides effective gradient updates to the shared weights in the network.
For the second stage, we stick to the simple data filtering $F_{\text{ADM}}(\mathcal{D}_{\text{ADM}},\mathcal{Y}_{k})=(\mathcal{D}_{\text{real},\mathcal{Y}_{k}},\mathcal{C}_{\mathcal{Y}_{k}})$, so that both adversarial and KL losses focus on the corresponding partition, given that each student has enough mode coverage from the first stage.
4.3 Distilling smaller students from scratch
Via the frameworks presented in the last two sections, MSD enables a performance upgrade over alternatives for one student with the same model architecture. In this section, we investigate training multiple students with smaller architectures – and thus faster inference time – without impairing much performance. However, this requires distilling into a student with a different architecture, preventing initialization from pretrained teacher weights. Distilling a single-step student from scratch has previously been difficult [70], and we could not obtain competitive results with the simple pipeline in Eq. 7. Therefore, we propose an additional pretraining phase $\text{Distill}_{\text{TSM}}$, with TSM denoting Teacher Score Matching, to find a good initialization for single-step distillation. TSM employs the following score-matching loss:
$\mathcal{L}_{\text{TSM}}=\mathbb{E}_{t}[\lambda_{t}\|\bm{\mu}_{\text{TSM}}^{\varphi}(\bm{x}_{t},t)-\bm{\mu}_{\text{teacher}}(\bm{x}_{t},t)\|_{2}^{2}],$ (8)
where the smaller student with weights $\varphi$ is trained to match the teacher’s score on real images at different noise levels. This step provides useful initialization weights for single-step distillation and is crucial to ensure convergence. With TSM added, the whole pipeline now becomes:
$\bm{\mu}^{(0)}$ $=\text{Distill}_{\text{TSM}}\left(\bm{\mu}_{\text{teacher}},\mathcal{D}_{\text{real}}\right),$ (9)
$G_{k}^{(1)}$ $=\text{Distill}_{\text{DM}}\left(\bm{\mu}_{\text{teacher}},F_{\text{DM}}(\mathcal{D}_{\text{DM}},\mathcal{Y}_{k});\bm{\mu}^{(0)}\right),\ \ k=1,...,K,$
$G_{k}^{(2)}$ $=\text{Distill}_{\text{ADM}}\left(\bm{\mu}_{\text{teacher}},F_{\text{ADM}}(\mathcal{D}_{\text{ADM}},\mathcal{Y}_{k});G_{k}^{(1)}\right),\ \ k=1,...,K.$
Although a smaller student may not perfectly match the teacher’s score, it still provides a good initialization for stages 1 and 2. The performance gap is remedied in the latter stages by focusing on a smaller partition for each student. This three-stage training scheme is illustrated in Fig. 3.
5 Experiments
To evaluate the effectiveness of our approach, we trained MSD with different design choices and compared against competing methods, including other single-step distillation methods.
In Sec. 5.1, we compare single vs multiple students on a 2D toy problem for direct visual comparison. For these experiments, we used only the DM stage. In Sec. 5.2, we investigate class-conditional image generation on ImageNet-64$\times$64 [11] where we have naturally defined classes to partition. Here, we explored training with the DM stage only, with both DM and ADM stages, and with all three stages for smaller students. We then evaluate MSD for a larger model in Sec. 5.3. We explored text-to-image generation on MS-COCO2014 [34] with varying training stages. We use the standard Fréchet Inception Distance (FID) [18] score to measure generation quality. Comprehensive comparisons confirm that MSD outperforms single-student counterparts and achieves state-of-the-art performance in single-step diffusion distillation. Finally, in Sec. 5.4, we summarize our ablation experiments over design choices.
To focus on the performance boost from multi-student distillation, we applied minimal changes to the hyperparameters used by Yin et al. [76, 75] for their distribution matching distillation implementations. More details on training and evaluation can be found in the App. D and E.
[…]
Figure 4: A 2D toy model. From left to right: teacher (multi-step) generation and student, one-step generation with $1$ and $8$ distilled students, the $\ell_{1}$ distance of generated samples between teacher and students. Takeaway: More students improve distillation quality on this easy-to-visualize setup.
5.1 Toy Experiments
In Fig. 4, we show the sample density of MSD with DM stage for a 2D toy experiment, where the real data distribution has $8$ classes, and each class is a mixture of $8$ Gaussians. We used a simple MLP with EDM schedules to train the teacher and distill into $1$, $2$, $4$, and $8$ students for comparison. The displayed samples and the $\ell_{1}$ distance from teacher generation show the collective generation quality increases as the number of students increases. App. A.2 additionally shows the results of Consistency Distillation [65] in this setting, demonstrating our framework’s generality.
5.2 Class-conditional Image Generation
Student architecture the same as the teacher: We trained $K=4$ students using the MSD framework and the EDM [25] teacher on class-conditional ImageNet-64$\times$64 generation. We applied the simplest strategy for splitting classes among students: Each student is responsible for $250$ consecutive classes in numerical order (i.e., $\nicefrac{{1}}{{K}}$ of the $1000$ classes). We compare performance with prior methods and show results in Tab. 2. Our DM stage, which uses the complementary regression loss, surpasses the one-student counterpart DMD [76], achieving a modest $0.25$ drop in FID score, making it a strong competitor in single-step distillation without an adversarial loss. We then took the best pretrained checkpoints and trained with the ADM stage, again outperforming the single-student counterpart DMD2 [75] with an FID score of $1.20$. It surpasses even the EDM teacher, StyleGAN-XL [57], the multi-step RIN [23] due to the adversarial loss. Fig. 5(a) and (b) compare sample generations, showing our best students have comparable generation quality as the teacher.
Table 1: Comparing class-conditional generators on ImageNet-64$\times$64. The number of function evaluations (NFE) for MSD is $1$ as a single student is used at inference for the given input.
Method NFE ($\downarrow$) FID ($\downarrow$)
Multiple Steps
RIN [23] $1000$ $1.23$
ADM [12] $250$ $2.07$
DPM Solver [39] $10$ $7.93$
Multistep CD [17] $2$ $2.0$
Single Step, w/o GAN
PD [55] $1$ $15.39$
DSNO [83] $1$ $7.83$
Diff-Instruct [43] $1$ $5.57$
iCT-deep [63] $1$ $3.25$
Moment Matching [56] $1$ $3.0$
DMD [76] $1$ $2.62$
MSD (ours): 4 students, DM only $1$ $2.37$
EMD [70] $1$ $2.20$
SiD [88] $1$ $1.52$
Single Step, w/ GAN
Post-distillation, 4, 42% smaller students $1$ $11.67$
MSD (ours): 4, 42% smaller students, ADM $1$ $2.88$
StyleGAN-XL [57] $1$ $1.52$
CTM [27] $1$ $1.92$
DMD2 [75] $1$ $1.28$
MSD (ours): 4 students, ADM $1$ $1.20$
SiDA [87] $1$ $1.11$
teacher
EDM (teacher, ODE) [25] $511$ $2.32$
EDM (teacher, SDE) [25] $511$ $1.36$
Table 2: Comparing MSD to other methods on zero-shot text-to-image generation on MS-COCO2014. We measure speed with sampling time per prompt (latency) and quality with FID.
Method Latency ($\downarrow$) FID ($\downarrow$)
Unaccelerated
[…]
LDM [54] $3.7$s $12.63$
eDiff-I [3] $32.0$s $6.95$
GANs
StyleGAN-T [58] $0.10$s $12.90$
GigaGAN [77] $0.13$s $9.09$
Accelerated
DPM++ (4 step) [40] $0.26$s $22.36$
InstaFlow-0.9B [37] $0.09$s $12.10$
UFOGen [72] $0.09$s $12.78$
DMD [76] $0.09$s $11.49$
EMD [70] $0.09$s $9.66$
DMD2 (w/o GAN) $0.09$s $9.28$
MSD (ours): 4 students, DM only $0.09$s $8.80$
DMD2 [75] $0.09$s $8.35$
MSD (ours): 4 students, ADM $0.09$s $8.20$
SiD-LSG [86] $0.09$s $8.15$
teacher
SDv1.5 ($50$ step, CFG=3, ODE) $2.59$s $8.59$
SDv1.5 ($200$ step, CFG=2, SDE) $10.25$s $7.21$
Student architecture smaller than the teacher: Next, we trained $4$ smaller student models with the prepended teacher score matching (TSM) stage from Sec. 4.3. This achieved a $42\%$ reduction in model size and a $7\%$ reduction in latency, with a slight degradation in FID score, offering a flexible framework to increase generation speed by reducing student size and increasing generation quality by training more students. Fig. 5(c) displays sample generation from these smaller students, whereas Fig. 5(d) shows sample generations from an even smaller set of students, with a $71\%$ percent reduction in model size and a $23\%$ percent reduction in latency. We observed slightly degraded but still competitive generation qualities. Using more and larger students will further boost performance, as shown by ablations in Sec. 5.4 and App. B.4. Smaller students without the TSM stage fail to reach even proper convergence. Moreover, instead of the TSM stage, we performed post-output distillation on the best single-step checkpoints and observed a significant drop in performance. Hence, the TSM stage is both necessary and efficient.
Refer to captionRefer to captionRefer to captionRefer to caption(a) Teacher (multistep)(b) Same-sized students(c) 42% smaller students(d) 71% smaller students
Figure 5: Sample generations on ImageNet-64$\times$64 from the teacher and different sized students, with architecture and latency details in App. D. The same-size students have comparable or slightly better generation quality than the teacher. Smaller students achieve faster generation while still having decent qualities. Same-sized students are trained with DM and ADM stages, whereas smaller students are trained with all three stages as shown in Fig. 3.
5.3 Text-to-Image Generation
Student architecture the same as the teacher: We evaluated the performance of text-to-image generation using the MS-COCO2014 [34] evaluation dataset. We distilled $4$ students from Stable Diffusion (SD) v1.5 [54] on a 5M-image subset of the COYO dataset [8]. For splitting prompts among students, we again employed a minimalist design: pass the prompts through the pre-trained SD v1.5 text encoder, pool the embeddings over the temporal dimension, and divide into $4$ disjoint subsets along $4$ quadrants. We trained with a classifier-free guidance (CFG) scale of $1.75$ for best FID performance. Tab. 2 compares the evaluation results with previous methods. Our baseline method with only the DM stage again achieved a performance boost with a $0.48$ drop in FID over the single-student counterpart DMD2 without adversarial loss [75]. Continuing the ADM stage from the best checkpoint yields a terminal FID of $8.20$, surpassing the single-student counterpart and achieving a close-to state-of-the-art FID score. In addition, for better visual quality, we also train with a larger CFG scale of $8$ with corresponding samples in Fig. 2(b) and App. F.2.
Student architecture smaller than the teacher: As a preliminary exploration, with the prepended teacher score matching (TSM) stage, we train a $83\%$ smaller and $5\%$ faster student on a dog-related prompt subset of COYO (containing 1 210 000similar-toabsent1210000$\sim$\$1\,210\,000\$∼ 1 210 000 prompts). We trained with a CFG scale of $8$ and display the samples in Fig. 2. We observed fair generation quality despite a significant drop in model size. Improved training is likely to obtain better sample quality and generalization power. Due to limited computational resources and the complete coverage of the prompt set by the 4-student model, we did not train the full set of students at this size.
5.4 Ablation Studies
Table 3: Ablation studies on different components of MSD. All experiments are done on ImageNet-64$\times$64, trained with only the DM stage for $20$k iterations, where $B$ is the batch size per student. See App. B.1.
Method FID ($\downarrow$)
$4$ students, $B=32$ $2.53$
$1$ students, $B=128$ $2.60$
$2$ students, $B=128$ $2.49$
$\bf{4}$ students, $\bf{B=128}$ (baseline) $2.37$
$8$ students, $B=128$ $2.32$
$4$ students, $B=128$, $K$-means splitting $2.39$
$4$ students, $B=128$, random splitting $2.45$
Here, we ablate the effect of different components in MSD and offer insight into scaling. Unless otherwise mentioned, all experiments are conducted for class-conditional generation ImageNet-64$\times$64, using only the DM stage for computational efficiency.
MSD is still better with the same effective batch size. To investigate if the performance boost from MSD comes from only a batch size increase over single student distillation, we make a comparison with the same effective batch size. As showcased in Tab. 3, MSD with $4$ students and a batch size of $32$ per student performs slightly better than the single-student counterpart with a batch size of $128$, indicating that MSD likely benefits from a capacity increase than a batch size increase. As a takeaway, with a fixed training resource measured in processed data points, users are better off distilling into multiple students with partitioned resources each than using all resources to distill into a single student. Our previous experiments also reflected this, where we used significantly fewer resources per student than the single-student counterparts (see details in App. D). Although multiple students means multiple model weights to save, storage is often cheap, so in many applications, this cost is outweighed by our improved quality or latency.
Simple splitting works surprisingly well. We used consecutive splitting of classes in Sec.  5.2. Although it shows an obvious advantage over random splitting, as shown in Tab. 3, it does not use the embedding information from the pretrained EDM model. Therefore, we investigated another strategy where we performed a $K$-means clustering ($K=4$) on the label embeddings, resulting in $4$ clusters of similar sizes: $(230,283,280,207)$. However, MSD trained with these clustered partitions performs similarly to sequential partition, as shown in Tab. 3. For text-to-image generation, we performed $K$-means clustering on the pooled embeddings of prompts in the training data, resulting in clusters of vastly uneven sizes. Due to computational limitations, we opted for the simpler partition strategies outlined in Sec. 5.
Effect of scaling the number of students. In Tab. 3, we study the effect of increasing $K$, the number of distilled students. We fixed the per-student batch size so that more students would induce a larger effective batch size. We observe better FID scores for more students. We hypothesize that better training strategies, such as per-student tuning, will further improve the quality. Optimal strategies for scaling to ultra-large numbers of students is an interesting area for future work.
6 Discussion
6.1 Limitations
MSD is the first work to explore diffusion distillation with multiple students, and it admits a few limitations that call for future work. 1) Further explorations could offer more insights into optimal design choices for quality and latency on various datasets, such as the number of students, input condition size for each student, and other hyperparameters. This is especially beneficial if the training budget is limited. 2) We apply simple partitioning for both class- and text-conditions and assign them disjointly to different students. Although our empirical study shows that simple alternatives do not offer obvious advantages, more sophisticated routing mechanisms may help. 3) We use simple channel reduction when designing smaller students to demonstrate feasibility. This results in a significantly smaller latency reduction than the reduction in sample size. Exploring other designs of smaller students will likely increase their quality and throughput. 4) We train different students separately, but we expect that carefully designed weight-sharing, loss-sharing, or other interaction schemes can further enhance training efficiency. 5) We hypothesize that MSD can be applied to other state-of-the-art diffusion distillation methods and other modalities for similar benefits but leave this for future work.
6.2 Conclusion
This work presented Multi-Student Distillation, a simple yet efficient method to increase the effective model capacity for single-step diffusion distillation. We applied MSD to the distribution matching and adversarial distillation methods. We demonstrated their superior performance over single-student counterparts in class-conditional and text-to-image generations. Particularly, MSD with DMD2’s two-stage training achieves state-of-the-art FID scores. Moreover, we successfully distilled smaller students from scratch, demonstrating MSD’s potential in further reducing the generation latency with multiple smaller student distillations. We envision building on MSD to enable generation in real-time, enabling many new use cases.
Ethics Statements
Our work aims to improve the quality and speed of diffusion models. Thus, we may inherit ethics concerns from diffusion and generative models. Potential risks include fabricating facts or profiles that could mislead public opinion, displaying biased information that may amplify social biases, and displacing creative jobs from artists and designers.
Acknowledgements
We thank Amirmojtaba Sabour for their support with the 2D toy model and Marc Law for helpful discussions and feedback.
Appendix A Additional experimental results
A.1 CLIP score for high guidance scale
Table 5: CLIP score comparison for high guidance scale on MS-COCO2014. LCM-LoRA is trained with a classifier-free guidance (CFG) scale of $7.5$, while other methods use a CFG of $8$.
Method Latency ($\downarrow$) CLIP score ($\uparrow$)
DPM++ (4 step) [40] $0.26$s $0.309$
UniPC (4 step) [80] $0.26$s $0.308$
LCM-LoRA (1 step) [42] $0.09$s $0.238$
LCM-LoRA (4 step) [42] $0.19$s $0.297$
DMD2 (our reimplementation) [75] $0.09$s $0.306$
MSD4-ADM (ours) $0.09$s $0.308$
DMD [76] $0.09$s $0.320$
SDv1.5 (teacher) [54] $2.59$s $0.322$
Tab. 5 shows the CLIP score of MSD and single-student methods. MSD4-ADM achieves a competitive CLIP score and beats the single student counterpart, DMD2. We believe the CLIP score can increase further if one trains on the LAION dataset [61] instead of the COYO dataset [8].
A.2 Consistency Distillation, Toy Experiments
To show the wider applicability of MSD, we apply another distillation method, Consistency Distillation [65], on the toy experiment setting in Sec. 5.1. Fig. 6 displays generated samples and the $\ell_{1}$ distance from teacher generation. While noting that consistency distillation achieves a weaker distillation of the teacher in general, we again observe better performance for more students. This indicates the generality of MSD.
[…]
Figure 6: A 2D toy model using consistency distillation. From left to right: teacher (multi-step) generation and student, one-step generation with $1$ and $8$ distilled students, the $\ell_{1}$ distance of generated samples between teacher and students. Takeaway: More students improve distillation quality on this easy-to-visualize setup with consistency distillation.
Appendix B Additional ablation studies
B.1 Training curves for Sec. 5.4
Refer to captionRefer to captionRefer to captionFIDTraining iterationsTraining iterationsTraining iterations
Figure 7: FID comparisons during training for ablations in Table 3.
Fig. 7 displays the training curves for the ablation studies shown in Tab. 3. The relative terminal performances are also reflected in the training process.
B.2 The effect of paired dataset size on DMD
In Sec. 4.2, we mentioned the special filtering strategy for MSD at DM stage: instead of partitioning the paired dataset for corresponding classes, we choose to keep the same complete dataset for each student. Fig. 8 demonstrates that the alternative strategy discourages mode coverage and leads to a worse terminal performance.
Refer to captionFIDTraining iterations
Figure 8: Comparison of paired dataset filtering strategies for MSD4-DM. Partitioning the paired dataset for each student discourages mode coverage, which results in worse terminal performance. In comparison, keeping the same paired dataset for each student achieves better performance without impairing the convergence speed.
B.3 Single-step distillation methods comparison
Table 6: Comparison of various aspects of single-step distillation methods.
Method on ImageNet Terminal FID Iterated Images
DMD [76] (our reimplementation) $2.54$ $\sim$ $130$M
DMD2 (w/o GAN) [75] $2.61$ $\sim$ $110$M
EMD [70] $2.20$ $\sim$ $600$M
SiD ($\alpha=1.0$) [88] $2.02$ $\sim$ $500$M
CTM [27] $>$$5$ $\sim$ $3$M
Table 6 justifies our choice of DMD/DMD2 as our first-stage training without adversarial loss. Competitor methods either need larger training data size (EMD, SiD) or have worse quality (CTM and other CM-based methods). DMD/DMD2, on the other hand, strike a good balance. DMD exhibits more stability for MSD on ImageNet, whereas DMD2 performs better for SD v1.5, which leads to our respective choices. As mentioned in App. D, we used a smaller-sized paired dataset (10 0001000010\,00010 000 images) than the original DMD paper (25 0002500025\,00025 000 images) for ImageNet, which significantly accelerated convergence without impairing the final performance. Moreover, as in Sec. 4.2, the same paired dataset can be used for all students, eliminating potential additional computation.
B.4 More results on distilling into smaller students
Refer to captionFID on first $10$ classes# Classes per student
Figure 9: The effect of student size, number of classes covered, and training stage on the performance of a single smaller student. XS, XXS, and B refer to model sizes, with precise definitions provided in Tab. 7. Also, see special evaluation details in App. E.
In Sec. 5.2, we trained MSD4-ADM on smaller students to demonstrate the tradeoff between generation quality and speed. Here, we make a more comprehensive ablation study on the interplay between student size, number of classes covered, and training stage, with results in Fig. 9. We observe that generation quality increases with student size and decreases with more classes covered. MSD offers great flexibility for users to make these choices based on computational resources, generation quality, and inference speed requirements.
Appendix C Deployment suggestions
Here, we discuss some MSD deployment options for practitioners.
A naive option for deployment is to use increased GPU memory to host all models simultaneously. However, this is impractical and wasteful, as only a single student model needs to be used for each user request. In settings with less GPU memory than all students’ sum memory requirement, we must swap student models on and off GPUs. This incurs extra latency. However, in the few-GPU many-users setting, there are already prominent latency issues, such as users needing to queue for usage. In few-user settings, resources are likely being taken offline to save cost, and thus, there is start-up latency for fresh requests, too. Therefore, we argue that the more interesting setting is in large distributed deployment.
For settings with more GPU memory than all students’ sum memory requirements, we can distribute the student models among a cluster of GPUs (as one would the teacher) and route each generation request to the appropriate student node. The routing layer is lightweight compared to the inference cost, so we pay little for it.
If the data has been partitioned uniformly according to user demand, the incoming requests will be distributed uniformly among the student nodes. Therefore, we achieve equal throughput compared to the teacher without more overall model storage. However, finding such a partition is challenging, and user demand may change over time. This leaves finding the optimal allocation of resources to the student nodes an open problem. In practice, we expect that a reduced student model size would lead to an overall reduction in storage requirements compared to the teacher alone.
Appendix D Implementation Details
D.1 Toy Experiments
The real dataset is a mixture of Gaussians. The radius for the outer circle is $0.5$, the radius for the $8$ inner circles is $0.1$, and the standard deviation for each Gaussian (smallest circle) is $0.005$. The teacher is trained with EDM noise schedule, where use $(\sigma_{\min},\sigma_{\max})=(0.002,80)$ and discretized the noise schedule into $1000$ steps. We train the teacher for 100 000100000100\,000100 000 iterations with AdamW [38] optimizer, setting the learning rate at 1e-4, weight decay to $0.01$, and beta parameters to $(0.9,0.999)$. For distillation, we first generated a dataset of $1000$ pairs, then used DMD [76] to train $1$, $2$, $4$, and $8$ students, respectively, all for 200 000200000200\,000200 000 iterations, with reduced learning rate at 1e-7. We only sample the first $750$ of the $1000$ steps for distillation.
Each subfigure of Fig. 4 is a histogram with $200$ bins on 100 000100000100\,000100 000 generated samples, using a custom colormap. The loss is the mean absolute difference of binned histogram values.
D.2 ImageNet
D.2.1 Same-sized students
Our ImageNet experiments setup closely follows DMD [76] and DMD2 [75] papers. We distill our one-step generators using the EDM [25] checkpoint “edm-imagenet-64x64-cond-adm”. We use $\sigma_{\min}=0.002$ and $\sigma_{\max}=80$ and discretize the noise schedules into $1000$ bins. The weight $w_{t}$ in Eq. (2) is set to $\frac{\sigma_{t}^{2}}{\alpha_{t}}\frac{CS}{\|\bm{\mu}_{\text{teacher}}(\bm{x}_{t},t)-\bm{x}\|_{1}}$ where $S$ is the number of spatial locations and $C$ is the number of channels, and the weight $\lambda_{t}$ in Eq. (4) is set to $(\sigma_{t}^{2}+0.5^{2})/(\sigma_{t}\cdot 0.5)^{2}$.
For the DM stage, we prepare a distillation dataset by generating 10 0001000010\,00010 000 noise-image pairs using the deterministic Heun sampler (with $S_{\text{churn}}=0$ over $256$ steps. We use the AdamW optimizer [38] with learning rate 2e-6, weight decay $0.01$, and beta parameters $(0.9,0.999)$. We compute the LPIPS loss using a VGG backbone from the LPIPS library [79], and we upscale the image to $224\times 224$ using bilinear upsampling. The regression loss weight is set to $0.25$. We use mixed-precision training and a gradient clipping with an $\ell_{2}$ norm of $10$. We partition the $1000$ total classes into consecutive blocks of $250$ classes and trained $4$ specialized students using $\text{Distill}_{\text{DM}}{}$ and $F_{\text{DM}}{}$ defined in Sec. 7. Each student is trained on $4$ A100 GPUs, with a total batch size of $128$, for 200 000200000200\,000200 000 iterations. This yields the MSD4-DM checkpoint in Tab. 2.
For the ADM stage, we attach a prediction head to the middle block of the fakescore model. The prediction head consists of a stack of $4\times 4$ convolutions with a stride of $2$, group normalization, and SiLU activations. All feature maps are downsampled to $4\times 4$ resolution, followed by a single convolutional layer with a kernel size and stride of $4$. The final output linear layer maps the given vector to a scalar predicted probability. We load the best generator checkpoint from DM stage but re-initialize the fakescore and GAN classifier model from teacher weights, as we observed this leads to slightly better performance. We set the GAN generator loss weight to 3e-3 and the GAN discriminator loss weight to 1e-2 and reduce the learning rate to 5e-7. Each student is trained on $4$ A100 GPUs, with a total batch size of $192$, for 150 000150000150\,000150 000 iterations. This yields the MSD4-ADM checkpoint in Tab. 2.
D.2.2 Smaller students
Following a minimalist design, we pick our smaller student’s architecture by changing hyperparameter values of the “edm-imagenet-64x64-cond-adm” checkpoint architecture. See details in Tab. 7.
Table 7: Hyperparameter details for different sized student models of the ADM architecture. Unspecified hyperparameters remain the same as the teacher. Latency is measured on a single NVIDIA RTX 4090 GPU.
Model Identifier # Channels Channel Multipliers # Residual Blocks # Parameters Latency
B (teacher) $192$ $[1,2,3,4]$ $3$ $296$M $0.0271$s
S $160$ $[1,2,2,4]$ $3$ $173$M $0.0253$s
XS $128$ $[1,2,2,4]$ $2$ $86$M $0.0209$s
XXS $96$ $[1,2,2,2]$ $2$ $26$M $0.0192$s
For the MSD4-ADM-S checkpoint in Tab. 2, we train the TSM stage using the model architecture S with the continuous EDM noise schedule with $(P_{\text{mean}},P_{\text{std}})=(-1.2,1.2)$ and the weighting $\lambda_{t}=(\sigma_{t}^{2}+0.5^{2})/(\sigma_{t}\cdot 0.5)^{2}$. We use a learning rate of 1e-4. Each student is trained on $4$ A100 GPUs, with a total batch size of $576$, for 400 000400000400\,000400 000 iterations. Then DM and ADM stages were trained using a total batch size of $160$, following otherwise the same setup as Sec. D.2.1.
For the ablation study in B.4, we instead train a common TSM stage for all students for computational efficiency. We train this common stage using $16$ A100 GPUs, with a total batch size of $3584$ and $4864$ for architecture XS and XXS, respectively. The DM and ADM stages are followed by specialized students with filtered data and $4$ A100 GPUs each, with a total batch size of $224$ and $256$ for architecture XS and XXS, respectively, and using the same learning rate of 2e-6.
D.3 SD v1.5
D.3.1 Same-sized students, CFG=1.75
Our SD v1.5 experiments setup closely follows DMD2 [75] paper. We distill our one-step generators from the SD v1.5 [54] model, using a classifier-free guidance scale of $1.75$ for the teacher model to obtain the best FID score. We use the first $5$M prompts from the COYO dataset [8] and the corresponding $5$M images for the GAN discriminator. We apply the DDIM noise schedule with $1000$ steps for sampling $t$. The weight $w_{t}$ in Eq. 2 is set to $\frac{\sigma_{t}^{2}}{\alpha_{t}}\frac{CS}{\|\bm{\mu}_{\text{teacher}}(\bm{x}_{t},t)-\bm{x}\|_{1}}$ where $S$ is the number of spatial locations and $C$ is the number of channels, and the weight $\lambda_{t}$ in Eq. 4 is set to $\alpha_{t}^{2}/\sigma_{t}^{2}$.
For the DM stage, we use the AdamW optimizer [38] with learning rate 1e-5, weight decay $0.01$, and beta parameters $(0.9,0.999)$. We use gradient checkpointing, mixed-precision training, and a gradient clipping with an $\ell_{2}$ norm of $10$. We partition the prompts and corresponding images by the $4$ quadrants formed by the first $2$ entries of the embeddings, where the embeddings are pooled from the outputs of the SD v1.5 text embedding layers. We choose not to include a regression loss but instead use a TTUR, which updates the fakescore model $10$ times per generator update. Each of the $4$ students is trained on $32$ A100 GPUs, with a total batch size of $1536$, for 40 0004000040\,00040 000 iterations. This yields the MSD4-DM checkpoint in Tab. 2.
For the ADM stage, we attach a prediction head that has the same architecture (though with a different input size) as Sec. D.2. We load both the best generator checkpoint and the corresponding fakescore checkpoint from DM stage. We set the GAN generator loss weight to 1e-3 and the GAN discriminator loss weight to 1e-2 and reduced the learning rate to 5e-7. Each student is trained on $32$ A100 GPUs, with a total batch size of $1024$, for $5000$ iterations. This yields the MSD4-ADM checkpoint in Tab. 2.
D.3.2 Same-sized students, CFG=8
The above CFG $=1.75$ setting yields sub-optimal image quality. Similar to previous works [76, 33, 54], we choose CFG $=8$ for enhanced image quality. Due to time and computational resource limitations, we only train with the ADM stage. Each of the $4$ students is trained on $32$ A100 GPUs, with a learning rate of 1e-5 and a batch size of $1024$ for both the fake and the real images, for $6000$ iterations. This yields the checkpoint that is used to generate Fig. 2 (b) and Fig .12. Longer training with the added DM stage can likely further improve the generation quality.
D.3.3 Smaller student, CFG=8
We again pick our smaller student’s architecture by changing the hyperparameter values of the SD v1.5 architecture. See details in Tab. 8.
Table 8: Hyperparameter details for different sized student models of the SD v1.5 architecture. Only the diffusion model part is measured since the text encoder and the VAE remain frozen. Unspecified hyperparameters remain the same as the teacher. Latency is measured on a single NVIDIA RTX 4090 GPU.
Model Identifier # block_out_channels # Parameters Latency
B (teacher) $[320,640,1280,1280]$ $860$M $0.041$s
S $[160,320,320,640]$ $142$M $0.039$s
To create a subset of dog-related data, we first selected 1 210 000similar-toabsent1210000$\sim$\$1\,210\,000\$∼ 1 210 000 prompts in the COYO [8] dataset whose embeddings are closest to “a dog.” We then created an equal number of noise-image pairs from the SD v1.5 teacher using these prompts with CFG $=8$. We train the TSM stage using the model architecture S with the $1000$-step DDIM noise schedule and the weighting $\lambda_{t}=\alpha_{t}^{2}/\sigma_{t}^{2}$. We use a learning rate of 1e-4. We then continue to the DM stage with the paired regression loss, using a learning rate of 1e-5, and finally continue to the ADM stage using generated paired images as “real” images with a learning rate of 5e-7. We use $16$ A100 GPUs. The TSM stage is trained with a total batch size of $1536$ for 240 000240000240\,000240 000 iterations. The DM stage is trained with a total batch size of $512$ for both the paired and the fake images for 20 0002000020\,00020 000 iterations, and the ADM stage is trained with the same batch size for $6000$ iterations. This yields the checkpoint used to generate Fig. 2(c). Longer training and better tuning are again likely to improve the generation quality further.
Appendix E Evaluation Details
For zero-shot COCO evaluation, we use the exact setup as GigaGAN [77] and DMD2 [75]. Specifically, we generate 30 0003000030\,00030 000 images using the prompts provided by DMD2 code. We downsample the generated images using PIL to $256\times 256$ Lanczos resizer. We then use the clean-fid [49] to compute the FID score between generated images and 40 5044050440\,50440 504 real images from the COCO 2014 validation dataset. Additionally, we use the OpenCLIP-G backbone to compute the CLIP score. For ImageNet, we generate 50 0005000050\,00050 000 images and calculate FID using EDM [25] evaluation code. When selecting the best checkpoints for partitioned students, the same procedure is applied only for prompts/classes within respective partitions. For the ablation study in Sec. B.4, 10 0001000010\,00010 000 images are generated for only the first $10$ classes for an apple-to-apple comparison.
Appendix F Additional qualitative results
F.1 Additional Imagenet-64$\times$64 results
Figure 10: Collective one-step samples from $4$ same-sized students trained with MSD-ADM on ImageNet (FID $=1.20$).
Figure 11: Collective one-step samples from $4$ smaller students trained with MSD-ADM on ImageNet (FID $=2.88$).
In Fig. 10, we present more ImageNet-64$\times$64 qualitative results collectively generated by our $4$ same-sized students trained with MSD-ADM. In Fig. 11, we display corresponding generations from $4$ smaller students with architecture S (see Tab. 7).
F.2 Additional Text-to-image Synthesis Results
Figure 12: Collective one-step samples from $4$ SD v1.5 students trained with MSD-ADM on COYO with CFG $=8$.
In Fig. 12, we present more text-to-image qualitative results collectively generated by our $4$ students trained on the COYO dataset with MSD-ADM. These students are trained with a teacher classifier-free guidance (CFG) scale of $1.75$, which yields sub-optimal visual qualities despite having a good FID score. Generation with better qualities can be obtained with a higher CFG scale.
Appendix G Prompt details
G.1 Prompts for Fig. 12
We use the following prompts for Fig. 12, from left to right, top to bottom:
  • wise old man with a white beard in the enchanted and magical forest
  • a high-resolution photo of an orange Porsche under sunshine
  • Astronaut on a camel on mars
  • a hot air balloon in shape of a heart. Grand Canyon
  • transparent vacation pod at dramatic scottish lochside, concept prototype, ultra clear plastic material, editorial style photograph
  • penguin standing on a sidewalk
  • border collie surfing a small wave, with a mountain on background
  • an underwater photo portrait of a beautiful fluffy white cat, hair floating. In a dynamic swimming pose. The sun rays filters through the water. High-angle shot. Shot on Fujifilm X
  • 3D animation cinematic style young caveman kid, in its natural environment
  • robot with human body form, robot pieces, knolling, top of view, ultra realistic
  • 3D render baby parrot, Chibi, adorable big eyes. In a garden with butterflies, greenery, lush, whimsical and soft, magical, octane render, fairy dust
  • a chimpanzee sitting on a wooden bench
  • a capybara made of voxels sitting in a field
  • a cat reading a newspaper
  • a squirrell driving a toy car
  • close-up photo of a unicorn in a forest, in a style of movie still
G.2 Prompts for Fig. 2
We use the following prompts (same for all three models), from left to right, top to bottom:
  • dog on a bed
  • Your Puppy Your Dog
  • Trained Happy Dog
  • Very handsome dog.